OpenStudy (anonymous):

Simplify the given expression. square root negative 6 open parentheses 2 plus square root of negative 8

OpenStudy (anonymous):

please write like this 1/x+4*2-3/ ... USE / * - + instead of words... I don't understand it at all :S

OpenStudy (anonymous):

can you rewrite it ?

OpenStudy (anonymous):

\[\sqrt{-6}\left( 2 + \sqrt{-8} \right)\]

OpenStudy (anonymous):

6i(2+8i) =12i+48i^2 =12i-48

OpenStudy (anonymous):

\[\LARGE \sqrt{-6}\cdot (2+\sqrt{-8})=i6\cdot(2+i8)=12i+48i^2\] since \[\LARGE i^2=-1 \] \[\LARGE -1\cdot 48+12i\]

OpenStudy (anonymous):

complex numbers.

OpenStudy (anonymous):

none of those are my options though

OpenStudy (anonymous):

what are your options? post them...

OpenStudy (anonymous):